A multiple choice quiz has4 questions. Each question has 5 possible​ answers, only one of which is correct. A student has not studied for the quiz and guesses randomly on each question. What is the probability that the student gets at least one correct​ answer? Show your​ calculation, and write your answer as either a reduced fraction or a decimal rounded to four places.
Answer #1

We have here, n = 4 and p = 1/5 = 0.2  

The probability that the student gets at least one correct​ answer

=1-0.4096

=0.5904